At 9:30, a car A started its journey and travelled at 65 km/h. At 11:00, car B started its journey from the same place and travelled on the same road as car A. If car B, travelling at a constant speed, took five hours to catch up with car A, find the speed of car B.
Answer by mananth(16949)

Car A 65 km/h
Car B x km/h
Car A 09:30
Car B 11:00
Difference in time= 01:30 => 1.50 hours
Car A will have covered 97.50 miles when Car B starts
catch up distance= 97.50 miles
catch up speed = x -65
catch up time = 5 hours
t=d/r
5=97.50/(x-65)
5x-325=97.50
5x=325+97.50
5x=422.50
/5
x=84.62 km/h-------------Car B